# The Euclidean distance degree of curves: from rational to line multiview varieties

### v1

@@ 2 ED degree of multiview varieties of a rational curve
Assume we are given a generic camera arrangement $\mathbf{C}=(C_{1},\dots,C_{n})$ of full rank $(h+1)\times(N+1)$ matrices $C_{i}$ with $N\geq 3$ and $h\geq 2$. Let $f:\mathbb{P}^{1}\to\mathbb{P}^{N}$ be as in (2.1) so that $Y=f(\mathbb{P}^{1})$ is degree $e$ and $Y$ is either smooth or has only nodal singularities.
Then $\mathbf{C}\mathbin{\vbox{\hbox{$\scriptscriptstyle\square$}}}f(\mathbb{P}^{1})$ is either a smooth curve or only has nodes as singularities.
@@
For $n=1$, this is a standard algebraic geometry fact. For instance, see [8, Exercise 3.34]. The case when $n>1$ immediately follows. ∎
@@
If $h\geq 2$ then the ED degree of $(\mathbf{C}\mathbin{\vbox{\hbox{$\scriptscriptstyle\square$}}}f(\mathbb{P}^{1}))_{\operatorname{Aff}}$ is seven. In Figure 1 (left) we plot $\mathbf{C}\mathbin{\vbox{\hbox{$\scriptscriptstyle\square$}}}f(\mathbb{P}^{1})$ in orange for a camera of size $3\times 4$ and the curve $Y$ in yellow.
$\diamond$
@@
for all $i,j\in[n],\ i\neq j$, then $S_{\infty}$ has $ne$ points. This holds by genericity of the camera arrangement: the degree $e$ polynomial $C_{i}^{(0)}\cdot f$ has a multiple root if and only if its $X$-discriminant vanishes, and $C_{i}^{(0)}\cdot f,C_{j}^{(0)}\cdot f$ share a root if and only if their resultant is nonvanishing [11, Section 1.1]. Therefore, if $f$ is a generically one-to-one map then $H_{\infty}\cap X$ has precisely $ne$ points.
@@
is a homogeneous degree $2ne$ polynomial in $s,t$. Because $\beta_{0}$ is generic, if $g_{\beta}$ has no common roots with $D$ then $g_{\beta}$ has $2ne$ distinct roots. On the other hand, if $g_{\beta}$ had a common root with $D$, then there would exist a unique $i=a$ such that $\left(C_{i}^{(0)}\cdot f\right)$ vanishes on that root. This implies the polynomial
@@
If $X_{\operatorname{Aff}}$ is smooth, an application of [20, Theorem 3.8] implies that
@@
Suppose $f$ is as in (2.1) with $N\geq 3$ and $h\geq 2$. Let $\mathcal{C}\subset\mathbb{P}^{(h+1)\times(N+1)-1}$ be an irreducible projective variety. If $C_{1},\dots,C_{n}$ are generic points in $\mathcal{C}$ and
@@
$\mathcal{C}:=\left\{\begin{bmatrix}c_{1}&c_{2}&0&0&0&0\\ 0&0&c_{3}&c_{4}&0&0\\ 0&0&0&0&c_{5}&c_{6}\end{bmatrix}\in\mathbb{K}^{3\times 6}:c_{i}\in\mathbb{K}\right\}.$
If $C$ is a generic element in $\mathcal{C}$, then $\operatorname{affEDdeg}(C_{1}\mathbin{\vbox{\hbox{$\scriptscriptstyle\square$}}}Y)=9$. We see that this disagrees with the numbers in (2.11).
@@
$\mathcal{C}=\left\{[c_{i,j}]\in\mathbb{K}^{3\times 4}:c_{1,1}=0\right\}.$
Let $C_{1}$ and $C_{2}$ be generic points of $\mathcal{C}$. If $Y$ is the rational normal curve in $\mathbb{P}^{3}$ parameterized by $t^{3}:st^{2}:s^{2}t:s^{3}]$, then
@@
Let $C$ be a general $3\times{(4+1)}$ camera matrix. Then $C\mathbin{\vbox{\hbox{$\scriptscriptstyle\square$}}}X$ is a degree $4$ curve in $\mathbb{P}^{2}$ with $\binom{4-1}{2}$ nodal singularities. The set of smooth points of $C\mathbin{\vbox{\hbox{$\scriptscriptstyle\square$}}}X\subset\mathbb{P}^{2}$ has Euler characteristic $\chi(\mathbb{P}^{1}\setminus\{2\cdot\binom{4-1}{2}\text{ points}\})=2-2\cdot\binom{4-1}{2}$. The ED degree of $(C\mathbin{\vbox{\hbox{$\scriptscriptstyle\square$}}}X)_{\operatorname{Aff}}$ is $3\cdot 4-2$. The projective ED degrees of $X$ and $C\mathbin{\vbox{\hbox{$\scriptscriptstyle\square$}}}X$ both equal $3\cdot 4-2$. As affine cones in ${{\mathbb{C}}}^{4+1}$ and ${{\mathbb{C}}}^{3}$, these varieties are parameterized by degree $4$ polynomials and one might predict from [6, Section 2] that the ED degree is $(2\cdot 4-1)^{2}$. The reason we do not get this bound is because these are general homogeneous polynomials of degree $4$ and not general degree $4$ polynomials. $\diamond$
@@
The results in this section have an intersection theoretic [10] interpretation through the multigrees presented in Section 1.2. Namely, the proof in Theorem 2.3 is showing that the curve $\mathbf{C}\mathbin{\vbox{\hbox{$\scriptscriptstyle\square$}}}f(\mathbb{P}^{1})\subset(\mathbb{P}^{h})^{n}$ has multidegree $\sum_{i=1}^{n}e\cdot\frac{T_{1}^{h}T_{2}^{h}\cdots T_{n}^{h}}{T_{i}}$ and intersects transversally each of the hyperplanes $V(x_{i,0})$ for $i=1,\dots,n$. This implies $\chi(X\cap H_{\infty})=en$.
@@
(obtained by homogenizing the polynomial in (2.6)) also transversally intersects the curve $\mathbf{C}\mathbin{\vbox{\hbox{$\scriptscriptstyle\square$}}}f(\mathbb{P}^{1})$ with no solutions at infinity. Since the multidegree of $V(q_{\beta})$ is $\sum_{i=1}^{n}2T_{i}^{2}$, we have $\chi(X\cap V(q_{\beta}))=2en$ because
$\left(\sum_{i=1}^{n}2T_{i}^{2}\right)\cdot\left(\sum_{i=1}^{n}e\cdot\frac{T_{1}^{h}\cdots T_{n}^{h}}{T_{i}}\right)=2en\cdot T_{1}^{h}\cdots T_{n}^{h}\in\mathbb{Z}[T_{1},\dots,T_{n}]/\langle T_{1}^{h+1},\dots,T_{n}^{h+1}\rangle.$
3 Strengthening the ED-equivalence catalogue
In this section, we present our main application of our results. Section 3.1 defines our main varieties of interest: a set of multiview varieties anchored at a subvariety of the Grassmannian. Section 3.2 gives a complementary approach for obtaining these multiview varieties using wedge cameras. It also provides the necessary exterior algebra background. Using our results in Section 2, we prove a pair of conjectures in Section 3.3.
@@ 3.1 Anchoring the $n$-view variety
Section 1.5 considers the point multiview variety $\mathbf{C}\mathbin{\vbox{\hbox{$\scriptscriptstyle\square$}}}Y$ anchored at $Y$ as the Zariski closure of the image of the map (1.5) for $Y\subset\mathbb{P}^{N}$. In this section, we carry this idea over to a linear map between Grassmannians. The associated multiview variety (called a generalized multiview variety in [7]) is obtained as the Zariski closure of a rational map
@@
In [7], the full data of a multiview variety is recorded in the notation $\mathcal{M}_{\mathbf{C},k}^{\Lambda,N,h}$. We opt for the more parsimonious notation and refer to the catalogue in [7, Theorem 3.3] for a complete description of multiview varieties arising from projections of points and lines in 1, 2, and 3-dimensional projective space.
@@
h=2,n=2h=2,\,n=2 $\mathbb{P}^{2}\times\mathbb{P}^{2}$ $2T_{1}+2T_{2}$ 10
h=3,n=1h=3,\,n=1 $\mathbb{P}^{5}$ $2T_{1}$ 4
h=3,n=2h=3,\,n=2 $\mathbb{P}^{5}\times\mathbb{P}^{5}$ $2T_{1}+2T_{2}$ 10
@@ 3.2 Exterior algebra review and wedge cameras
Next, we show how to realize $\mathbf{C}^{h}\mathbin{\vbox{\hbox{$\scriptscriptstyle\square$}}}{\texttt{L}^{\texttt{3}}}$ as a point multiview variety anchored at $\iota({\texttt{L}^{\texttt{3}}})\subset\mathbb{P}^{5}$. To that end, we review the exterior algebra (see [18, Chapter XVI] for a textbook reference) to state Section 3.2.
@@
where $\bigwedge^{k}V$ is the space of alternating $k$-linear forms on $V$. Elements of $\bigwedge^{k}V$ are called $k$-vectors, and the wedge product satisfies $u\wedge v=-v\wedge u$ and $v\wedge v=0$. A simple wedge is one of the form $v_{1}\wedge\cdots\wedge v_{k}$, $v_{i}\in V.$ Every element of $\bigwedge^{k}V$ is a linear combination of simple wedges. The dimension of $\bigwedge^{k}V$ is $\binom{\dim V}{k}$, and a basis is given by
@@
${\wedge^{2}C}=\begin{bmatrix}-1&0&0&3&4&0\\ 0&5&0&10&0&-20\\ 0&0&0&-5&0&0\end{bmatrix}$.
@@ 3.3 Resolving a pair of ED degree conjectures
Let $\mathbf{D}^{h}=(D_{1},\dots,D_{n})$ denote the $2$-wedge camera configuration of $\mathbf{C}^{h}$.
@@
Since $\mathbf{C}^{h}$ is generic, each $D_{i}$ is a generic point in the irreducible variety of wedge cameras of size $\binom{h+1}{2}\times 6$. It follows from Figure 4 that $\iota_{1,3,1}({\texttt{L}^{\texttt{3}}})$ is parameterized by $f$ in Section 1.6. Thus, by Section 2, it suffices to show (3.5) for $n=1,2$ for $h=2,3$. This holds by Table 1 and the result follows. ∎
Remark 3.9 (Rational scrolls as curves in $\operatorname{Gr}(1,\mathbb{P}^{N})$).
The ideas used in the proof above can be put into a broader framework. In this remark, we provide an additional example to motivate the study of ED degrees of subvarieties of the Grassmannian.
@@
A Bezier curve in $\mathbb{R}^{N}$ of degree $E$ is defined by a set of control points $P_{0},\dots,P_{E}\in\mathbb{K}^{N}$. The curve is given parametrically by
$\mathbf{B}(s,t)=\sum_{i=0}^{E}B_{i,E}(s,t)\cdot{P}_{i},\quad s=1,\,t\in[0,1]$
@@
where $B_{i,E}(s,t)$ are the homogenized Bernstein polynomials $B_{i,E}(s,t)=\binom{E}{i}(s-t)^{E-i}t^{i}.$ This parameterization has the advantage that the curve segment $\{\mathbf{B}(1,t):t\in[0:1]\}$ on the standard affine chart in $\mathbb{R}^{N}$ is in the convex hull of the control points, starts at $P_{0}$, and ends at $P_{E}$. It follows from Theorem 2.3 that if the control points are in generic position, then $\operatorname{affEDdeg}(\mathbf{C}\mathbin{\vbox{\hbox{$\scriptscriptstyle\square$}}}\mathbf{B}(s,t))=En-2$ for a generic camera configuration $\mathbf{C}$.
Suppose now we have a pair of Bezier curves $\mathbf{B}_{i}(1,t)$, $i=1,2$, in $\mathbb{K}^{3}$ with degree $E_{i}$. We obtain a one-dimensional family of lines by considering the line spanned by $\mathbf{B}_{1}(1,t)$ and $\mathbf{B}_{2}(1,t)$ for $t\in[0,1]$. This is visualized in Figure 5. Again we assume the control points are in generic position. The Zariski closure of the family of lines is a subvariety of $\operatorname{Gr}(1,\mathbb{P}^{3})$, and we denote it as $\Lambda_{\mathbf{B}_{1},\mathbf{B}_{2}}$. The Plücker embedding of this variety is given by the two by two minors of the $2\times 4$ matrix
@@
This gives a degree $E_{1}+E_{2}$ polynomial map from $\mathbb{P}^{1}$ to $\mathbb{P}^{5}$, and the results of Theorem 2.3 imply
$\operatorname{affEDdeg}(\mathbf{C}\mathbin{\vbox{\hbox{$\scriptscriptstyle\square$}}}\Lambda_{\mathbf{B}_{1},\mathbf{B}_{2}})\,=\,3(E_{1}+E_{2})n-2$
where $\mathbf{C}$ is an arrangement of $n$ cameras of size $3\times 4$.
Figure 5: The purple ruled surfaces consist of line segments connecting a pair of Bezier curves of degree one (left) and four (right). The lines spanned by the segments in a ruling give $\Lambda_{\mathbf{B}_{1},\mathbf{B}_{2}}\subset\operatorname{Gr}(1,\mathbb{P}^{3})$. The left surface has two rulings as discussed in Section 1.6.

### v2

@@ 2 ED degrees of multiview varieties anchored at rational curves
Assume we are given a generic camera arrangement $\mathbf{C}=(C_{1},\dots,C_{n})$ of full rank $(h+1)\times(N+1)$ matrices $C_{i}$ with $N\geq 3$ and $h\geq 2$. Let $f:\mathbb{P}^{1}\to\mathbb{P}^{N}$ be as in (2.1) so that $Y=f(\mathbb{P}^{1})$ is degree $E$ and $Y$ is either smooth or has only nodal singularities. Then $\mathbf{C}\mathbin{\vbox{\hbox{$\scriptscriptstyle\square$}}}f(\mathbb{P}^{1})$ is either a smooth curve or only has nodes as singularities.
@@
For $n=1$, this is a standard algebraic geometry fact. For instance, see [10, Exercise 3.34]. The case when $n>1$ immediately follows. ∎
@@
If $h\geq 2$ then the ED degree of $(\mathbf{C}\mathbin{\vbox{\hbox{$\scriptscriptstyle\square$}}}f(\mathbb{P}^{1}))_{\operatorname{Aff}}$ is seven. In Figure 1 (left) we plot $\mathbf{C}\mathbin{\vbox{\hbox{$\scriptscriptstyle\square$}}}f(\mathbb{P}^{1})$ in orange for a camera of size $3\times 4$ and the curve $Y$ in yellow. $\diamond$
@@
for all $i,j\in[n],\ i\neq j$, then $S_{\infty}$ has $En$ points. This holds by genericity of the camera arrangement: the degree $E$ polynomial $C_{i}^{(0)}\cdot f$ has a multiple root if and only if its $X$-discriminant vanishes, and $C_{i}^{(0)}\cdot f,C_{j}^{(0)}\cdot f$ share a root if and only if their resultant is nonvanishing [14, Section 1.1]. Therefore, if $f$ is a generically one-to-one map then $H_{\infty}\cap X$ has precisely $En$ points.
@@
is a homogeneous degree $2En$ polynomial in $s,t$. Because $\beta_{0}$ is generic, if $g_{\beta}$ has no common roots with $D$ then $g_{\beta}$ has $2En$ distinct roots. On the other hand, if $g_{\beta}$ had a common root with $D$, then there would exist a unique $i=a$ such that $\left(C_{i}^{(0)}\cdot f\right)$ vanishes on that root. This implies the polynomial
@@
If $X_{\operatorname{Aff}}$ is smooth, an application of [27, Theorem 3.8] implies that
@@
Suppose $f$ is as in (2.1) with $N\geq 3$ and $h\geq 2$. Let $\mathcal{A}\subset\mathbb{P}^{(h+1)\times(N+1)-1}$ be an irreducible projective variety. If $C_{1},\dots,C_{n}$ are generic points in $\mathcal{A}$ and
@@
$\mathcal{A}:=\left\{\begin{bmatrix}c_{1}&c_{2}&0&0&0&0\\ 0&0&c_{3}&c_{4}&0&0\\ 0&0&0&0&c_{5}&c_{6}\end{bmatrix}\in\mathbb{K}^{3\times 6}:c_{i}\in\mathbb{K}\right\}.$
If $C$ is a generic element in $\mathcal{A}$, then $\operatorname{affEDdeg}(C_{1}\mathbin{\vbox{\hbox{$\scriptscriptstyle\square$}}}Y)=9$. We see that this disagrees with the numbers in (2.11).
@@
$\mathcal{A}=\left\{[c_{i,j}]\in{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbb{P}^{3\times 4-1}}:c_{1,1}=0\right\}.$
Let $C_{1}$ and $C_{2}$ be generic points of $\mathcal{A}$. If $Y$ is the rational normal curve in $\mathbb{P}^{3}$ parameterized by $t^{3}:st^{2}:s^{2}t:s^{3}]$, then
@@
Let $C$ be a general $3\times{(4+1)}$ camera matrix. Then $C\mathbin{\vbox{\hbox{$\scriptscriptstyle\square$}}}X$ is a degree $4$ curve in $\mathbb{P}^{2}$ with $\binom{4-1}{2}$ nodal singularities. The set of smooth points of $C\mathbin{\vbox{\hbox{$\scriptscriptstyle\square$}}}X\subset\mathbb{P}^{2}$ has Euler characteristic $\chi(\mathbb{P}^{1}\setminus\{2\cdot\binom{4-1}{2}\text{ points}\})=2-2\cdot\binom{4-1}{2}$. The ED degree of $(C\mathbin{\vbox{\hbox{$\scriptscriptstyle\square$}}}X)_{\operatorname{Aff}}$ is $3\cdot 4-2$. The projective ED degrees of $X$ and $C\mathbin{\vbox{\hbox{$\scriptscriptstyle\square$}}}X$ both equal $3\cdot 4-2$. As affine cones in ${{\mathbb{C}}}^{4+1}$ and ${{\mathbb{C}}}^{3}$, these varieties are parameterized by degree $4$ polynomials and one might predict from [8, Section 2] that the ED degree is $(2\cdot 4-1)^{2}$. The reason we do not get this bound is because these are general homogeneous polynomials of degree $4$ and not general degree $4$ polynomials. $\diamond$
@@
The results in this section have an intersection theoretic [13] interpretation through the multidegrees presented in Section 1.2. Namely, the proof in Theorem 2.3 is showing that the curve $\mathbf{C}\mathbin{\vbox{\hbox{$\scriptscriptstyle\square$}}}f(\mathbb{P}^{1})\subset(\mathbb{P}^{h})^{n}$ has multidegree $\sum_{i=1}^{n}E\cdot{T_{1}^{h-1}T_{2}^{h-1}\cdots T_{n}^{h-1}}$ and intersects transversally each of the hyperplanes $V(x_{i,0})$ for $i=1,\dots,n$. This implies $\chi(X\cap H_{\infty})=En$.
@@
(obtained by homogenizing the polynomial in (2.6)) also transversally intersects the curve $\mathbf{C}\mathbin{\vbox{\hbox{$\scriptscriptstyle\square$}}}f(\mathbb{P}^{1})$ with no solutions at infinity. Since the multidegree of $V(q_{\beta})$ is $\sum_{i=1}^{n}2T_{i}$, we have $\chi(X\cap V(q_{\beta}))=2En$ because
$\left(\sum_{i=1}^{n}2T_{i}\right)\cdot\left(\sum_{i=1}^{n}E\cdot{T_{1}^{h-1}\cdots T_{n}^{h-1}}\right)=2En\cdot T_{1}^{h}\cdots T_{n}^{h}\in\mathbb{Z}[T_{1},\dots,T_{n}]/\langle T_{1}^{h+1},\dots,T_{n}^{h+1}\rangle.$
3 Settling conjectures of Duff–Rydell
In this section, we present our main application of our Theorem 2.3. Section 3.1 defines our main varieties of interest: a set of multiview varieties anchored at a subvariety of the Grassmannian. Section 3.2 gives a complementary approach for obtaining these multiview varieties using wedge cameras. It also provides the necessary exterior algebra background. Using our results in Section 2, we prove a pair of conjectures in Section 3.3.
@@ 3.1 Anchoring the $n$-view variety
Section 1.5 considers the point multiview variety $\mathbf{C}\mathbin{\vbox{\hbox{$\scriptscriptstyle\square$}}}Y$ anchored at $Y$ as the Zariski closure of the image of the map (1.6) for $Y\subset\mathbb{P}^{N}$. In this section, we carry this idea over to a map between Grassmannians. The associated multiview variety (called a generalized multiview variety in [9]) is obtained as the Zariski closure of a rational map
@@
In [9], the full data of a multiview variety is recorded in the notation $\mathcal{M}_{\mathbf{C},k}^{\Lambda,N,h}$. We opt for the more parsimonious notation and refer to the catalogue in [9, Theorem 3.3] for a complete description of multiview varieties arising from projections of points and lines in 1, 2, and 3-dimensional projective space.
@@
h=2,n=2h=2,\,n=2 $\mathbb{P}^{2}\times\mathbb{P}^{2}$ $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(2T_{1}+2T_{2})T_{1}T_{2}$} 10
h=3,n=1h=3,\,n=1 $\mathbb{P}^{5}$ $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}2T_{1}^{4}$} 4
h=3,n=2h=3,\,n=2 $\mathbb{P}^{5}\times\mathbb{P}^{5}$ $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(2T_{1}+2T_{2})T_{1}^{4}T_{2}^{4}$} 10
@@ 3.2 Exterior algebra review and wedge cameras
Next, we show how to realize $\mathbf{C}^{h}\mathbin{\vbox{\hbox{$\scriptscriptstyle\square$}}}{\texttt{L}^{\texttt{3}}}$ as a point multiview variety anchored at $\iota({\texttt{L}^{\texttt{3}}})\subset\mathbb{P}^{5}$. To that end, we review the exterior algebra (see [25, Chapter XVI] for a textbook reference) to state Section 3.2.
@@
where $\bigwedge^{k}V$ is the space of $k$-vectors, the dual of the space of alternating $k$-linear forms on $V$. The wedge product satisfies $u\wedge v=-v\wedge u$ and $v\wedge v=0$. A simple wedge is one of the form $v_{1}\wedge\cdots\wedge v_{k}$, $v_{i}\in V.$ Every element of $\bigwedge^{k}V$ is a linear combination of simple wedges. The dimension of $\bigwedge^{k}V$ is $\binom{\dim V}{k}$, and a basis is given by
@@
${\wedge^{2}C}=\begin{bmatrix}-1&0&3&0&4&0\\ 0&5&10&0&0&-20\\ 0&0&-5&0&0&0\end{bmatrix}$.
@@ 3.3 Resolving a pair of ED degree conjectures
Let $h=2$ or $h=3$, and let $\mathbf{D}^{h}=(D_{1},\dots,D_{n})$ denote the $2$-wedge camera configuration of $\mathbf{C}^{h}$.
@@
Since $\mathbf{C}^{h}$ is generic, each $D_{i}$ is a generic point in the irreducible variety of wedge cameras of size $\binom{h+1}{2}\times 6$. It follows from Figure 4 that $\iota_{1,3,1}({\texttt{L}^{\texttt{3}}})$ is parameterized by $f$ in Section 1.6. Thus, by Section 2, it suffices to show (3.5) for $n=1,2$. This holds by Table 1 and the result follows. ∎
4 One-parameter families of 3D lines
We study an additional model to further motivate the investigation of ED Degrees of subvarieties of the Grassmannian. One important example comes from rational scrolls. These are surfaces that represent one-parameter families of 3D lines and arise naturally in computer vision applications. In this section, we consider an instance of such a family: lines sweeping along Bézier curves to provide a one-parameter line-view variety. This is obtained by mapping a subvariety of $\operatorname{Gr}(1,\mathbb{P}^{3})$ to $\operatorname{Gr}(1,\mathbb{P}^{2})^{n}$.
@@ 4 One-parameter families of 3D lines
For background (e.g., [15, Chapter 5]), recall a Bézier curve in $\mathbb{K}^{N}$ of degree $E$ is defined by a set of $E+1$ control points
$P_{0},\dots,P_{E}\in\mathbb{K}^{N}$
with $N\geq 2$. The curve is given parametrically by
$\mathbf{B}(1,t)=\sum_{i=0}^{E}B_{i,E}(1,t)\cdot{P}_{i},\quad\,t\in[0,1]$ (4.1)
where $B_{i,E}(s,t)$ is the $i$th homogenized Bernstein polynomial of degree $E$,
$B_{i,E}(s,t):=\binom{E}{i}(s-t)^{E-i}t^{i}.$
@@
This parameterization has the advantage that the curve segment $\{\mathbf{B}(1,t):t\in[0,1]\}$ on the standard affine chart in $\mathbb{K}^{N}$ is in the convex hull of the control points, starts at $P_{0}=\mathbf{B}(1,0)$, and ends at $P_{E}=\mathbf{B}(1,1)$. Using the homogenized polynomials, the parameterization from (4.1) extends to a projective map
$\mathbb{P}^{1}\longrightarrow\mathbb{P}^{N},\qquad[s:t]\longmapsto[s^{E}:\mathbf{B}(s,t)].$
We now construct a one-parameter family of lines. For simplicity, we assume $N=3$. Suppose we have a pair of Bézier curves $\mathbf{B}_{1}(1,t)$ and $\mathbf{B}_{2}(1,t)$ in $\mathbb{K}^{3}$ with degree $E_{1}$ and $E_{2}$, respectively. Their projective closures are the rational curves
$X_{i}:=\left\{[s^{E_{i}}:\mathbf{B}_{i}(s,t)]\in\mathbb{P}^{3}:[s:t]\in\mathbb{P}^{1}\right\}.$
Under genericity assumptions of the control points, we may assume $X_{1}\cap X_{2}$ is empty. Thus, the points
$s^{E_{1}}:\mathbf{B}_{1}(s,t)]\text{ and }[s^{E_{2}}:\mathbf{B}_{2}(s,t)]$
are distinct for all $s:t]\in\mathbb{P}^{1}$, and span a line in $\mathbb{P}^{3}$. This set of lines yields a subvariety of $\operatorname{Gr}(1,\mathbb{P}^{3})$, which we denote by $\Lambda_{\mathbf{B}_{1},\mathbf{B}_{2}}$. This is visualized in Figure 5.
Let $\mathcal{A}\subset\mathbb{P}^{11}$ be an irreducible projective variety of $3\times 4$ camera matrices. If $C_{1},\dots,C_{n}$ are generic points in $\mathcal{A}$, our line-multiview variety of interest is
$(C_{1},\dots,C_{n})\mathbin{\vbox{\hbox{$\scriptscriptstyle\square$}}}\Lambda_{\mathbf{B}_{1},\mathbf{B}_{2}}\;\subset\;\operatorname{Gr}(1,\mathbb{P}^{2})^{n}.$
Recalling Section 3.1, we say it is anchored at $\Lambda_{\mathbf{B}_{1},\mathbf{B}_{2}}$ with respect to the camera arrangement $\mathbf{C}=(C_{1},\dots,C_{n})$.
Theorem 4.1.
Let $\mathcal{A}\subset\mathbb{P}^{3\times 4-1}$ be an irreducible projective variety, and fix generic points $C_{1},\dots,C_{n}$ in $\mathcal{A}$. For Bézier curves $\mathbf{B}_{1}$ and $\mathbf{B}_{2}$ with $E_{1}+1$ and $E_{2}+1$ generic control points, let $X_{\mathbf{B}_{1},\mathbf{B}_{2},n}$ denote the multiprojective variety
$\iota\left(\,(C_{1},\dots,C_{n})\mathbin{\vbox{\hbox{$\scriptscriptstyle\square$}}}\Lambda_{\mathbf{B}_{1},\mathbf{B}_{2}}\,\right)\subset(\mathbb{P}^{2})^{n}.$
If
$\operatorname{affEDdeg}(X_{\mathbf{B}_{1},\mathbf{B}_{2},n})=3(E_{1}+E_{2})n-2,$ (4.2)
holds for $n=1,2$, then (4.2) holds for all $n\geq 1$.
Proof.
Under the Plücker embedding a line in $\Lambda_{\mathbf{B}_{1},\mathbf{B}_{2}}$ is determined by the six maximal minors of the $2\times 4$ matrix
@@
This yields a degree $E_{1}+E_{2}$ rational map $f_{\mathbf{B}_{1},\mathbf{B}_{2}}:\mathbb{P}^{1}\to\mathbb{P}^{5}$, assuming the genericity of the control points of the Bézier curves.
By the commutative diagram in Section 3.2, which is a consequence of Section 3.2, we have
$X_{\mathbf{B}_{1},\mathbf{B}_{2},n}={(D_{1},\dots,D_{n})\mathbin{\vbox{\hbox{$\scriptscriptstyle\square$}}}\iota(\Lambda_{\mathbf{B}_{1},\mathbf{B}_{2}})}$
where $D_{i}=\wedge^{2}C_{i}$. Since ${\iota(\Lambda_{\mathbf{B}_{1},\mathbf{B}_{2}})}=f_{\mathbf{B}_{1},\mathbf{B}_{2}}(\mathbb{P}^{1})$ is a rational curve in $\mathbb{P}^{3}$ of degree $E_{1}+E_{2}$, the result follows from Section 2 as the $D_{i}$ are generic points in the irreducible variety $\{\wedge^{2}C:C\in\mathcal{A}\}\subset\mathbb{P}^{3\times 6-1}$.
Figure 5: The purple ruled surfaces consist of line segments connecting a pair of Bézier curves of degree one (left) and four (right). The lines spanned by the segments in a ruling give $\Lambda_{\mathbf{B}_{1},\mathbf{B}_{2}}\subset\operatorname{Gr}(1,\mathbb{P}^{3})$. The left surface has two rulings as discussed in Section 1.6.